# Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning

### v1

@@ Abstract
Traditionally, learning the structure of a Dynamic Bayesian Network has been centralized, with all data pooled in one location. However, in real-world scenarios, data are often dispersed among multiple parties (e.g., companies, devices) that aim to collaboratively learn a Dynamic Bayesian Network while preserving their data privacy and security. In this study, we introduce a federated learning approach for estimating the structure of a Dynamic Bayesian Network from data distributed horizontally across different parties. We propose a distributed structure learning method that leverages continuous optimization so that only model parameters are exchanged during optimization. Experimental results on synthetic and real datasets reveal that our method outperforms other state-of-the-art techniques, particularly when there are many clients with limited individual sample sizes.
@@ 1 Introduction
The learning of Dynamic Bayesian Network (DBN) structures is a crucial technique in causal learning and representation learning. More importantly, it has been applied in various real-world applications, such as modeling gene expression [21], measuring network security [8], studying manufacturing process parameters [47], and recognizing interaction activities [54]. Traditionally, the learning of DBNs has been conducted at a centralized location where all data is aggregated. However, with the rapid advancements in technology and the Internet of Things (IoT) [27] devices, data collection has become significantly more accessible. Consequently, in real-world scenarios, data is typically owned and dispersed among multiple entities, including mobile devices, individuals, companies, and healthcare institutions.
In many cases, these entities, referred to as clients, may lack an adequate number of samples to independently construct a meaningful DBN. They might seek to collaboratively gain comprehensive insights into the DBN structure without disclosing their individual data due to privacy and security concerns. For instance, consider several hospitals aiming to work together to develop a DBN that uncovers the conditional independence relationships among their variables of interest, such as various medical conditions. Sharing patients’ records is clearly prohibited by privacy regulations [2]. Consequently, the sensitive nature of the data renders the centralized aggregation of data from different clients for DBN learning infeasible.
A robust privacy-preserving learning strategy is therefore essential. In recent years, federated learning (FL) [31] has gained significant attention, enabling numerous clients to collaboratively train a machine learning model through a coordinated and privacy-preserving approach. In this framework, each client is not required to share its raw data; instead, they only provide the minimal necessary information, such as model parameters and gradient updates, pertinent to the specific learning task. This approach has been applied across various research domains, including the Internet of Things [20] and recommender systems [7]. Readers may refer to these review papers for more comprehensive insights [23, 51].
@@
Traditionally, score-based learning of DBNs relies on discrete optimization, making it incompatible with common continuous optimization approaches used in federated learning methods. Fortunately, recent work by Zheng et al. [55] and Pamfil et al. [36] have made continuous optimization strategies possible for DBN learning by providing algebraic characterizations of acyclicity. Despite this advancement, directly applying popular federated optimization techniques (FedAvg [31], Ditto [22]) to these continuous formulations remains challenging due to the acyclicity constraint inherent in DBN structure learning.
@@
In this work, we present a federated learning approach for estimating the structure of Dynamic Bayesian Network from observational data (with temporal information) that is horizontally partitioned across different agents. Our contributions are threefold.
  • We propose a Federated DBN Learning method 2Dbn based on continuous optimization using the Alternating Direction Method of Multipliers (ADMM). 2Dbn can simultaneously learn contemporaneous and time-lagged causal structure without any implicit assumptions on the underlying graph topology.
@@
  • Using extensive simulations and two real-world, high-dimensional datasets, we demonstrate that 2Dbn is scalable and effectively identifies the true underlying DBN graph.
@@
  • To our knowledge, this is the first study to explore learning DBN within the FL framework.
@@
Our full code is available on GitHub: https://github.com/PeChen123/2DBN_learning.
Literature Review.
@@
Dynamic Bayesian Network are probabilistic models that extend static Bayesian Networks to represent multi-stage processes by modeling sequences of variables over different stages. Learning DBN involves two key tasks: structure learning (identifying the network topology) and parameter learning (estimating the conditional probabilities). Friedman et al. [12] laid the foundation for this field by developing methods to learn the structure of dynamic probabilistic networks, effectively extending Bayesian Networks to capture temporal dynamics in sequential data. Building on this work, Murphy [32] adapted static Bayesian Network learning techniques for temporal domains, employing methods such as the Expectation-Maximization algorithm for parameter estimation. Advancements in structure learning from time series data were further propelled by Pamfil et al. [36], who developed an extension of Bayesian Network inference using continuous optimization methods. Their approach enabled efficient and scalable learning of DBN structures by formulating the problem as a continuous optimization task with acyclicity constraints tailored for temporal graphs. To address non-linear dependencies in time series data, Tank et al. [48] introduced neural Granger causality methods, which leverage neural networks within the DBN framework to capture complex non-linear interactions. In the field of bioinformatics, Yu et al. [52] enhanced Bayesian Network inference techniques to generate causal networks from observational biological data. Their work demonstrated the practical applicability of DBN in the investigation of intricate biological relationships and regulatory mechanisms. However, traditional DBN learning methods often require centralized data aggregation, which can pose significant privacy risks and scalability challenges, particularly in sensitive domains such as healthcare and genomics. These limitations have spurred interest in privacy-preserving and distributed learning frameworks. Using Federated Learning, DBNs can be trained collaboratively across distributed datasets, enabling secure and scalable model inference while preserving data privacy.
@@
To the best of our knowledge, there is very limited work on federated or privacy-preserving approaches for learning Dynamic Bayesian Networks. However, there are numerous studies based on distributed learning for Bayesian Networks. Gou et al. [16] adopted a two-step procedure that first estimates the BN structures independently using each client’s local dataset and then applies a further conditional independence test. Na and Yang [33] proposed a voting mechanism to select edges that were identified by more than half of the clients. These approaches rely solely on the final graphs that are independently derived from each local dataset, which might result in suboptimal performance due to the limited exchange of information. More importantly, Ng and Zhang [35] proposed a distributed Bayesian learning method based on continuous optimization, using ADMM. Notably, no prior work has tackled distributed learning of DBN.
@@ 2 Problem Statement
We consider a scenario with a total of $K$ clients, indexed by $k\in\{1,2,\ldots,K\}$. Each client $k$ possesses its own local dataset consisting of $M$ realizations of a stationary time series. Specifically, for the $k$-th client, we have time series $\{x_{m,t}^{k}\}_{t=0}^{T}$ for $x_{m,t}^{k}\in\mathbb{R}^{d}$, where $m\in\{1,2,\ldots,M\}$ indexes each realization, and $d$ is the number of variables. Due to privacy regulations, these time series cannot be shared directly among clients. For clarity and conciseness, the main text omits the subscript $m$ corresponding to multiple time series, with the complete $m$ notation detailed in the supplementary material §7.1, and consider a generic realization $\{x_{t}^{k}\}_{t=0}^{T}$. In this setting, we assume that the data from all clients follow the same underlying distribution. The extension to non-homogeneous distributions is left for future work (see Sec. 6).
For each client $k$, its dataset (for a single realization) can be written as $\mathcal{D}^{k}=\{x_{t}^{k}\}_{t=p}^{T}$, where $t\in\{p,p+1,\ldots,T\}$ and $p$ denotes the autoregressive order. We model our data using Structural Equation Modeling (SEM) and consider a Vector Autoregressive structure (SVAR) of order $p$:
$(x_{t}^{k})^{\top}=(x_{t}^{k})^{\top}W+(x_{t-1}^{k})^{\top}A_{1}+\dots+(x_{t-p}^{k})^{\top}A_{p}+(u_{t}^{k})^{\top},$ (1)
@@
    $u_{t}^{k}\sim\mathcal{N}(0,I)$ is a vector of centered noise terms, independent across time and variables;
@@
    $W$ is a weighted adjacency matrix capturing intra-slice dependencies (edges at the same time step);
@@
    $A_{i}$, for $i\in\{1,\dots,p\}$, are the coefficient matrices capturing inter-slice dependencies (edges across different time steps).
For each client $k$, the effective sample size is $n_{k}=T+1-p$. Given the collection of datasets $X=\bigcup_{k=1}^{K}\mathcal{D}^{k}$ from all clients, the objective is to recover the true $W$ and $A$ in a privacy-preserving manner. We focus on a setup where clients are motivated to collaborate on learning the DBN structure but are only willing to share minimal information, such as model parameters or estimated graphs, about their local datasets. This setup involves horizontal partitioning of the dataset, which means that, while all local datasets share the same set of variables, the sample data differ between clients. The total sample size, $n$, is the sum of the local sample sizes: $n=\sum_{k=1}^{K}n_{k}$.
3 Federated DBN Inference with ADMM
To implement Federated learning to DBN, our method builds on the DYNOTEARS approach introduced by Pamfil et al. [36], which frames the structural learning of linear Dynamical Bayesian Networks as a continuous constrained optimization problem. By incorporating $\ell_{1}$-norm penalties to encourage sparsity, the optimization problem is formulated as follows:
$\begin{split}\min_{W,A}&\ \ell(W,A)\quad\text{subject to}\quad W\text{ is acyclic},\\ &\ell(W,A)=\frac{1}{2n}\|X_{t}-X_{t}W-X_{(t-p):(t-1)}A\|_{F}^{2}+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}.\end{split}$ (2)
For the acyclicity constraint, following Zheng et al. [55],
@@ 3 Federated DBN Inference with ADMM
The ADMM [5] is an optimization algorithm designed to solve convex problems by breaking them into smaller, more manageable subproblems. It is particularly effective for large-scale optimization tasks with complex constraints. By following ADMM framework, we decompose the constrained problem (2) into multiple subproblems and utilize an iterative message-passing approach to converge to the final solution. ADMM proves particularly effective when subproblems have closed-form solutions, which we derive for the first subproblem. To cast problem (2) in an ADMM framework, we reformulate it using local variables $B_{1},\dots,B_{K}\in\mathbb{R}^{d\times d}$ for the intra-slice matrices and $D_{1},\dots,D_{K}\in\mathbb{R}^{d\times d}$ for the inter-slice matrices, along with shared global variables $W\in\mathbb{R}^{d\times d}$ and $A\in\mathbb{R}^{d\times d}$, representing an equivalent formulation:
@@
The local variables $B_{1},\dots,B_{K}$ and $D_{1},\dots,D_{K}$ represent the model parameters specific to each client. Notably, this problem resembles the global variable consensus ADMM framework. The constraints $B_{k}=W$ and $D_{k}=A$ are imposed to enforce consistency, ensuring that the local model parameters across clients remain identical.
Since ADMM combines elements of dual decomposition and the augmented Lagrangian method, making it efficient for handling separable objective functions and facilitating parallel computation, we employ the augmented Lagrangian method to transform the constrained problem into a series of unconstrained subproblems. The augmented Lagrangian is given by:
($\mathcal{L}\left(\{B_{k},D_{k}\}_{k=1}^{K},W,A,\alpha,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\rho_{1},\rho_{2}\right)=$
$\sum_{k=1}^{K}\Bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{\top}(B_{k}-W)\right)+\frac{\rho_{2}}{2}\|B_{k}-W\|_{F}^{2}$
@@
where $\{\beta_{k},\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$ and $\alpha\in\mathbb{R}$ are estimations of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are the penalty coefficients and $\|*\|_{F}$ is Frobenius norm. Then, we have the iterative update rules of ADMM as follows:
@@
where $\phi_{1},\phi_{2}\in\mathbb{R}$ are hyperparameters that control how fast the coefficients $\rho_{1},\rho_{2}$ are increased.
As previously mentioned, ADMM is particularly efficient when the optimization subproblems have closed-form solutions. The subproblem in equation (3) is a well-known proximal minimization problem, extensively explored in the field of numerical optimization [9, 37]. For readability, we define the following matrices:
@@
For a comprehensive procedure, we have summarized it in the supplementary material §7.4. The presence of the acyclicity term $h(W)$ prevents us from obtaining a closed-form solution for the problem in equation (4). Instead, the optimization problem can be addressed using first-order methods, such as gradient descent, or second-order methods, like L-BFGS [6]. In this work, we reproduce the approach of Pamfil et al. [36] and employ the L-BFGS method to solve the optimization problem. We have summarized our procedure in Algorithm 1.
Algorithm 1 Distributed DBN Inference with ADMM

### v2

# Federated Learning of Dynamic Bayesian Network via Continuous Optimization from Time Series Data

@@ Abstract
Traditionally, learning the structure of a Dynamic Bayesian Network has been centralized, requiring all data to be pooled in one location. However, in real-world scenarios, data are often distributed across multiple entities (e.g., companies, devices) that seek to collaboratively learn a Dynamic Bayesian Network while preserving data privacy and security. More importantly, due to the presence of diverse clients, the data may follow different distributions, resulting in data heterogeneity. This heterogeneity poses additional challenges for centralized approaches. In this study, we first introduce a federated learning approach for estimating the structure of a Dynamic Bayesian Network from homogeneous time series data that are horizontally distributed across different parties. We then extend this approach to heterogeneous time series data by incorporating a proximal operator as a regularization term in a personalized federated learning framework. To this end, we propose FDBNL and PFDBNL, which leverage continuous optimization, ensuring that only model parameters are exchanged during the optimization process. Experimental results on synthetic and real-world datasets demonstrate that our method outperforms state-of-the-art techniques, particularly in scenarios with many clients and limited individual sample sizes.
@@ 1 Introduction
Learning the structure of a Dynamic Bayesian Network (DBN) is a fundamental technique in causal learning and representation learning. DBNs are particularly valuable for uncovering temporal dependencies and conditional independence relationships in time-series data, making them a cornerstone of causal inference in dynamic systems. More importantly, DBNs have been widely applied in real-world domains, such as modeling gene expression [26], assessing network security [10], analyzing manufacturing process parameters [55], and recognizing interaction activities [62].
Traditionally, DBN structure learning has been conducted in a centralized setting, where all data is aggregated in one location. However, with the rapid advancements in technology and the proliferation of Internet of Things (IoT) [31] devices, data collection has become significantly more decentralized. In real-world scenarios, data is often owned and distributed across multiple entities, such as mobile devices, individuals, companies, and healthcare institutions. Many of these entities, referred to as clients, may lack a sufficient number of samples to independently construct a meaningful DBN. To gain comprehensive insights, they may seek to collaboratively learn the DBN structure while preserving data privacy and security. For instance, consider several hospitals aiming to collaboratively develop a DBN to uncover conditional independence relationships among medical conditions. Sharing patient records is strictly prohibited by privacy regulations [2], making centralized data aggregation infeasible. Furthermore, data heterogeneity introduces additional complexities: clients may exhibit homogeneous data (similar statistical distributions) or heterogeneous data, where distributions differ significantly or feature spaces (e.g., variables or attributes) vary entirely. These disparities complicate collaborative learning, as models must reconcile mismatched patterns or structural differences without direct data sharing.
A robust privacy-preserving learning strategy is therefore essential. In recent years, federated learning (FL) [35] has gained significant attention as a framework that enables multiple clients to collaboratively train machine learning models in a privacy-preserving manner. In FL, clients do not share raw data but instead exchange minimal necessary information, such as model parameters and gradient updates, relevant to the learning task. This approach has been successfully applied across various domains, including the Internet of Things [25], condition monitoring [4, 63], and recommender systems [8]. Readers seeking further insights into FL can refer to comprehensive review papers [28, 59]. Moreover, Personalized Federated Learning (PFL) extends FL to address the challenges posed by heterogeneous data distributions across clients. Unlike traditional FL, which trains a single global model, PFL allows for the development of personalized models tailored to each client’s unique data distribution while still benefiting from collaborative learning. To achieve this balance between personalization and collaboration, various PFL approaches have been proposed, including proximal regularization [48] to balance local and global models, clustered federated learning [16, 49, 36] to group clients with similar data distributions, and meta-learning-based methods [22, 9] to facilitate fast local adaptation. These techniques enhance performance in real-world federated environments while preserving data privacy and minimizing communication costs.
@@
Traditionally, score-based DBN learning relies on discrete optimization, making it incompatible with common continuous optimization techniques. However, continuous optimization methods, particularly those used in federated learning, solve problems efficiently using convex optimization techniques. Recent works by Zheng et al. [64] and Pamfil et al. [41] have addressed this limitation by introducing continuous optimization strategies for DBN learning, leveraging algebraic characterizations of acyclicity. Despite this advancement, directly applying popular federated optimization techniques - such as FedAvg [35] and Ditto [27] - to these continuous formulations remains challenging due to the inherent acyclicity constraint in DBN structure learning.
@@
In this work, we introduce a federated learning approach and a personalized federated learning approach for estimating the structure of a Dynamic Bayesian Network from homogeneous and heterogeneous observational time-series data that is horizontally partitioned across different clients. Our contributions can be summarized as follows:
  • Novel Federated Approach for DBN Learning:
    • Federated DBN Learning (FDBNL) for homogeneous time series, formulated as a continuous optimization problem and solved using the Alternating Direction Method of Multipliers (ADMM).
    • Personalized Federated DBN Learning (PFDBNL) for heterogeneous time series, incorporating a proximal operator to account for client-specific variations.
  • Both FDBNL and PFDBNL enable causal structure learning directly from data without imposing restrictive assumptions on the underlying graph topology.
@@
  • Through extensive simulations and experiments on two high-dimensional, real-world datasets, we demonstrate that our methods are scalable and effectively recover the true underlying DBN structures, even in challenging heterogeneous settings.
@@
  • To the best of our knowledge, this is the first study to explore DBN structure learning within the federated learning framework. Additionally, we pioneer the study of temporal dynamic structure learning from heterogeneous data in a federated context.
@@
Our full code is available on GitHub: https://github.com/PeChen123/FDBN_Learning.
Literature Review
@@
Dynamic Bayesian Networks are probabilistic models that extend static Bayesian Networks to represent multi-stage processes by modeling sequences of variables across different time steps. Learning a DBN involves two key tasks: structure learning (identifying the network topology) and parameter learning (estimating the conditional probabilities). Friedman et al. [15] laid the foundation for this field by developing methods to learn the structure of dynamic probabilistic networks, effectively extending Bayesian Networks to capture temporal dynamics in sequential data. Building on this work, Murphy [37] adapted static Bayesian Network learning techniques for temporal domains, employing methods such as the Expectation-Maximization algorithm for parameter estimation. Advancements in structure learning from time-series data were further propelled by Pamfil et al. [41], who developed an extension of Bayesian Network inference using continuous optimization methods. Their approach enabled efficient and scalable learning of DBN structures by formulating the problem as a continuous optimization task with acyclicity constraints tailored for temporal graphs. To address non-linear dependencies in time-series data, Tank et al. [56] introduced neural Granger causality methods, which leverage neural networks within the DBN framework to capture complex non-linear interactions. In the field of bioinformatics, Yu et al. [60] enhanced Bayesian Network inference techniques to generate causal networks from observational biological data. Their work demonstrated the practical applicability of DBNs in investigating intricate biological relationships and regulatory mechanisms. However, traditional DBN learning methods often require centralized data aggregation, which poses significant privacy risks and scalability challenges, particularly in sensitive domains such as healthcare and genomics. These limitations have spurred interest in privacy-preserving and distributed learning frameworks. Federated Learning enables DBNs to be trained collaboratively across distributed datasets, facilitating secure and scalable model inference while preserving data privacy.
@@
To the best of our knowledge, there is very limited work on federated or privacy-preserving approaches for learning Dynamic Bayesian Networks. However, there are numerous studies on distributed learning for Bayesian Networks. Gou et al. [20] adopted a two-step procedure that first estimates the BN structures independently using each client’s local dataset and then applies a further conditional independence test. Na and Yang [38] proposed a voting mechanism to select edges that were identified by more than half of the clients. These approaches rely solely on the final graphs independently derived from each local dataset, which may lead to suboptimal performance due to the limited exchange of information. More importantly, Ng and Zhang [40] proposed a distributed Bayesian Network learning method based on continuous optimization using the Alternating Direction Method of Multipliers (ADMM). Notably, no prior work has addressed distributed learning of DBNs.
@@ 2 Problem Statement
We consider a scenario with $K$ clients, indexed by $k\in\{1,2,\ldots,K\}$. Each client $k$ possesses a local dataset containing $M$ realizations of a stationary time series. Specifically, for the $k$-th client, the time series dataset is denoted as $\mathcal{D}^{k}=\{x_{m,t}^{k}\}_{t=0}^{T}$, where $x_{m,t}^{k}\in\mathbb{R}^{d}$, $m\in\{1,2,\ldots,M\}$ indexes realizations, and $d$ is the number of variables. Due to privacy constraints, these time series cannot be directly shared between clients. For simplicity, we omit the realization subscript $m$ in the main text (see supplementary material §7.1 for full notation) and consider a generic realization $\{x_{t}^{k}\}_{t=0}^{T}$.
For each client $k$, we model the data $\{x_{t}^{k}\}$ (for a single realization) using a Structural Equation Model (SEM) with a Structural Vector Autoregressive (SVAR) structure of order $p$:
$(x_{t}^{k})^{\top}=(x_{t}^{k})^{\top}W_{k}+(x_{t-1}^{k})^{\top}A_{k_{1}}+\dots+(x_{t-p}^{k})^{\top}A_{k_{p}}+(u_{t}^{k})^{\top},$
@@
    $t\in\{p,p+1,\ldots,T\}$, and $p$ is the autoregressive order.
@@
    $u_{t}^{k}\sim\mathcal{N}(0,I)$ is a noise vector, independent across time and variables.
@@
    $W_{k}$ is a weighted adjacency matrix (structured as a directed acyclic graph (DAG)) capturing intra-slice dependencies for the $k$-th client.
  4. 4.
    $A_{k_{i}}$ ($i\in\{1,\dots,p\}$) are coefficient matrices capturing inter-slice dependencies for the $k$-th client.
The matrices $W_{k}$ and $A_{k_{i}}$ may be homogeneous (shared across all clients) or heterogeneous (client-specific). We assume the order $p$ is same for all clients. In the homogeneous case, we write $W_{k}=W$ and $A_{k_{i}}=A_{i}$ since $W_{k}$ and $A_{k_{i}}$ are the same for all clients. In the heterogeneous case, these matrices vary by client. The sample size for client $k$ is $n_{k}=T+1-p$.
Given the federated dataset $X=\bigcup_{k=1}^{K}\mathcal{D}^{k}$, the goal is to recover the true parameters $W,A_{i}$ (homogeneous) or $W_{k},A_{k_{i}}$ (heterogeneous) in a privacy-preserving manner. Clients collaborate to learn the DBN structure, combining intra-slice ($W_{k}$) and inter-slice ($A_{k_{i}}$) dependencies, while sharing only minimal information (e.g., model parameters or graphs). The total sample size is $n=\sum_{k=1}^{K}n_{k}$. We also consider partial client participation, where only a subset of clients $j\leq K$ contributes in each communication round. This is critical for scalability, as inactive clients can rejoin later without disrupting the process. Notably, partial participation is primarily relevant in the heterogeneous case, where client-specific parameters require updates from participating clients. In the homogeneous case, shared parameters remain unaffected by non-participating clients.
3 Proposed Methodologies
In this section, we introduce the framework for Federated Dynamic Bayesian Network Learning and Personalized Federated Dynamic Bayesian Network Learning. The general idea is summarized in Figure 1.
Figure 1: Overview of Personalized Federated DBN learning with $d=3$ nodes, autoregression order $p=2$, and $K=3$ clients. The only difference between it and Federated DBN learning is that Federated DBN learning requires each $W_{k}$ and $A_{k_{i}}$ to be identical across clients.
3.1 Federated Dynamic Bayesian Network Learning
In this section, we introduce the Federated DBN learning (FDBNL) for the homogeneous (i.i.d) time series. To implement federated learning to DBN, our method builds on the method introduced by Pamfil et al. [41], which frames the structural learning of linear DBN as a continuous constrained optimization problem. By incorporating $\ell_{1}$-norm penalties to encourage sparsity, the optimization problem is formulated as follows:
$\begin{split}\min_{W,A}&\ \ell(W,A)+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}\quad\text{subject to}\quad W\text{ is acyclic},\\ &\ell(W,A)=\frac{1}{2n}\|X_{t}-X_{t}W-X_{(t-p):(t-1)}A\|_{F}^{2},\end{split}$ (1)
where $X_{(t-p):(t-1)}=\left[X_{t-1}\ \cdots\ X_{t-p}\right]$, with each $X_{t-i}$ (for $i\in\{1,\ldots,p\}$) being a time-lagged version of $X_{t}$ and $n\times d$ matrix where rows correspond to realizations. The matrix $A$ is defined as:
$A=\begin{bmatrix}A_{1}^{\top}\\ \vdots\\ A_{p}^{\top}\end{bmatrix},$
where each $A_{i}\in\mathbb{R}^{d\times d}$ captures lag-$i$ dependencies. Thus, we observe that $X_{(t-p):(t-1)}$ is an $n\times(p\cdot d)$ matrix, while $A$ is a $(p\cdot d)\times d$ matrix. For the acyclicity constraint, following Zheng et al. [64],
@@ 3.1 Federated Dynamic Bayesian Network Learning
This setting does not work for federated learning since all data is encapsulated. Thus, for the federated setting, we leverage the Alternating Direction Method of Multipliers [6], an optimization algorithm designed to solve convex problems by decomposing them into smaller, more manageable subproblems. It is particularly effective for large-scale optimization tasks with complex constraints. By following ADMM framework, we decompose the constrained problem (1) into multiple subproblems and utilize an iterative message-passing approach to converge to the final solution. ADMM proves particularly effective when subproblems have closed-form solutions, which we derive for the first subproblem. To cast problem (1) in an ADMM framework, we reformulate it using local variables $B_{1},\dots,B_{K}\in\mathbb{R}^{d\times d}$ for the intra-slice matrices and $D_{1},\dots,D_{K}\in\mathbb{R}^{pd\times d}$ for the inter-slice matrices, along with shared global variables $W\in\mathbb{R}^{d\times d}$ and $A\in\mathbb{R}^{pd\times d}$, representing an equivalent formulation:
@@
The local variables $B_{1},\dots,B_{K}$ and $D_{1},\dots,D_{K}$ represent the model parameters specific to each client. The data for the $k$-th client consists $X_{t}^{k}$ and $X^{k}_{(t-p):(t-1)}$. Notably, this problem resembles the global variable consensus ADMM framework. The constraints $B_{k}=W$ and $D_{k}=A$ are imposed to enforce consistency, ensuring that the local model parameters across clients remain identical. Since ADMM combines elements of dual decomposition and the augmented Lagrangian method, making it efficient for handling separable objective functions and facilitating parallel computation, we employ the augmented Lagrangian method to transform the constrained problem into a series of unconstrained subproblems. The augmented Lagrangian is given by:
($\mathcal{L}\left(\{B_{k},D_{k}\}_{k=1}^{K},W,A,\alpha,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\rho_{1},\rho_{2}\right)$ $=\sum_{k=1}^{K}\Bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{\top}(B_{k}-W)\right)+\frac{\rho_{2}}{2}\|B_{k}-W\|_{F}^{2}$
@@
where $\{\beta_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$, $\{\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{pd\times d}$ and $\alpha\in\mathbb{R}$ are estimations of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are the penalty coefficients and $\|\cdot\|_{F}$ is Frobenius norm. Then, we have the iterative update rules of ADMM as follows:
@@
where $\phi_{1},\phi_{2}\in\mathbb{R}$ are hyperparameters that control how fast the coefficients $\rho_{1},\rho_{2}$ are increased. As previously mentioned, ADMM is particularly efficient when the optimization subproblems have closed-form solutions. The subproblem in equation (2) is a well-known proximal minimization problem, extensively explored in the field of numerical optimization [11, 42]. For readability, we define the following matrices:
@@
For a comprehensive procedure, we have summarized it in the supplementary material §7.4. The presence of the acyclicity term $h(W)$ prevents us from obtaining a closed-form solution for the problem in equation (3). Instead, the optimization problem can be addressed using first-order methods, such as gradient descent, or second-order methods, like L-BFGS [7]. In this work, we write $W=W_{+}-W_{-}$ (and analogously for $A$) to handle the $\ell_{1}$ penalty term and employ the L-BFGS method to solve the optimization problem. We have summarized our procedure in Algorithm 1.
Algorithm 1 Federated Dynamic Bayesian Network Learning